# Just Ramp-up: Unleash the Potential of Regression-based Estimator for A/B Tests under Network Interference

### v1

@@ Abstract
A/B tests on modern platforms, including social networks and two-sided marketplaces, often encounter challenges due to network interference. Recent research in causal inference under network interference has explored various experimental designs and estimation techniques to address this issue. However, existing methods, which typically rely on single experiments, often reach a performance bottleneck and face limitations in handling diverse interference structures. In contrast, we propose leveraging multiple experiments to overcome these limitations. In industry, the use of sequential experiments, often known as the ramp-up process, where traffic to the treatment gradually increases, is common due to operational needs like risk management and cost control. Our approach shifts the focus from operational aspects to the statistical advantages of merging data from multiple experiments. By combining data from sequentially conducted experiments, we aim to estimate the global average treatment effect (GATE) more effectively. In this paper, we begin by analyzing the bias and variance of the linear regression estimator for GATE under general linear network interference. We demonstrate that bias plays a dominant role in the bias-variance tradeoff and highlight the intrinsic bias reduction achieved by merging data from experiments with strictly different treatment proportions. Herein the improvement introduced by merging two steps of experimental data is essential. In addition, we show that merging more steps of experimental data is unnecessary under general linear interference, while it can become beneficial when nonlinear interference occurs. We also provide an intuitive and informative explanation for this significant bias reduction, as well as the synergetic effect between our merging methodology and cluster-level randomization. Furthermore, we look into a more advanced estimator based on graph neural networks (GNN). Through extensive simulation studies [^1] The code is available at https://github.com/Cqyiiii/Just-Ramp-up., including novel and challenging scenarios such as square root interference, multihop interference, and dynamic graph structures, we show that the regression-based estimator benefits remarkably from training on merged experiment data, achieving outstanding statistical performance. Additionally, we also verify that most of our theoretical results and insights derived from linear regression can also be generalized to the GNN-based estimators.
@@ 1 Introduction
A/B tests, also known as randomized controlled trials, have been widely adopted by modern online platforms as a paradigm of data-driven decision-making to support rapid product iteration [30, 42]. In most cases, the interested causal estimand is the GATE, since the experimenter is concerned with how much improvement on performance metrics, such as user engagement, would be achieved by introducing a new product feature to all users versus maintaining the status quo. In complex systems like social networks and marketplaces, the classic Stable Unit Treatment Value Assumption (SUTVA) in causal inference is often violated in cases of social interaction and competition, where the outcome of one user is influenced by the treatments of neighboring users. In such scenarios, the challenge of causal inference under interference naturally arises. This interference, mediated through a graph—such as a social network or market bipartite—is called network interference.
@@ 2.1 Preliminary
We adopt the Neyman-Rubin’s potential outcome framework [59] and consider a finite population of $n$ units. The treatment vector is represented as $\mathbf{z}=\left(z_{1},z_{2},\ldots,z_{n}\right)\in\{0,1\}^{n}$. Without assuming the Stable Unit Treatment Value Assumption (SUTVA), we define the potential outcome of unit $i$ as $Y_{i}=Y_{i}(\mathbf{z})$ to account for its possible dependency on the treatments of other units.
@@ 2.2 General Linear Interference Model
Where $\beta_{0}$, $\beta_{1}$, and $\beta_{2}$ are unknown parameters, and we assume $\beta_{2}\neq 0$, indicating that there is indeed an interference effect. This model can be written more compactly as:
@@
If we take $B=D^{-1}A$, then our potential outcome model is reduced to the classic linear model, formatted as follows:
@@
This model is also called as linear-in-means model [2], which is widely analyzed and used for generating potential outcomes in simulation studies. We mention that removing the degree normalization, i.e. setting $B=A$ results in another common case where the exposure condition shifts from the proportion to the number of treated neighbors, thereby characterizing a different class of social interaction.
@@
$B=(\mathbf{1}\mathbf{1}^{T}-I_{n})\odot\left(\sum_{m=1}^{M}r_{m}(D^{-1}A)^{m}\right)$ (6)
@@ 2.4 Regression-based Estimator
$\hat{\tau}(f)=\frac{1}{n}\mathbf{1}^{T}\left(f(\mathbf{1},A)-f(\mathbf{0},A)\right)$ (8)
@@ 3 Analysis on Linear Regression Estimator
In this section, we perform a detailed analysis of the simple linear regression model [^2] We refer to it as ”simple” because it does not incorporate knowledge of graph topology; that is, $f(\mathbf{z},A)=f(\mathbf{z})$ for $f\in\mathcal{F}_{lr}$. As will be demonstrated in the subsequent subsections, this regression model provides a clear illustration of the role of merging experimental data for training. Many of the phenomena observed with this simple regression model can also be generalized to other regression models, such as those in $\mathcal{F}_{gnn}$, which will be discussed in the simulation study. In summary, the linear regression model examined here is both straightforward and sufficiently powerful for illustration, allowing for the development of insightful conclusions. To be specific, the considered linear regression model is given by:
$Y=X(\mathbf{z})\theta$ (9)
here $X(\mathbf{z})=(\mathbf{1},\mathbf{z})$, $\theta=(\theta_{0},\theta_{1})^{T}$.
@@
$\hat{\theta}=(X(\mathbf{z})^{T}X(\mathbf{z}))^{-1}X(\mathbf{z})^{T}Y$ (11)
@@ 3.1 One-step Experiment
$\operatorname{Bias}(\hat{\tau})$ $=\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B))$ (12)
$=\frac{\beta_{2}}{nc(1-c)}\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-c\mathbf{1})^{T}\right]B\right)-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)$
@@
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}$ $=(\frac{1}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}\right)+\frac{1}{\left(\beta_{2}nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{T}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (13)
@@
$\mathbf{1}^{T}D^{-1}A^{2}D^{-1}\mathbf{1}=O(n)$ (21)
For instance, in a ring graph, we have $\sum_{j}(\sum_{i}B_{ij})=\Theta(n)$, while in a star graph—where one opinion leader connects to all other units—we have $\sum_{j}(\sum_{i}B_{ij})=\Theta(n^{2})$. In the context of Example 1, a sufficient condition for satisfying this assumption is restricting the growth rate to a constant $\kappa$, a condition commonly employed in related works [64]. However, a restricted growth rate is not a necessary condition. As a counterexample, consider the complete graph, where $\mathbf{1}^{T}D^{-1}A^{2}D^{-1}\mathbf{1}=n$; despite being the densest graph, it lacks any prominent opinion leaders. We further assert that this assumption is reasonable for typical social networks like WeChat, where, even with over a billion nodes, it is unlikely for any single unit to have more than 100,000 friends while maintaining significant influence over each one of them.
@@ 3.3 $T$-step Experiment
Through Cauchy-Schwarz inequality and basic algebra, we can verify that
@@
In addition, we can apply the Corollary 2 in the reverse direction: given the proportions, $\{c_{1},...,c_{T}\}$, removing the data with treatment proportion $c_{t}$ from the merged data set could potentially reduce bias further. Repeating this procedure suggests the final solution would involve merging only the experimental datasets with proportions $c_{1}$ and $c_{T}$. This outcome underscores the particular significance of the $T=2$ case, and we will further verify it in the simulation study. Nevertheless, merging multistep experimental data can be beneficial under nonlinear interference, as discussed in Section 5. There, we will also outline the recommended practice for applying our methodology, in conjunction with the results of the simulation study.
@@
The simpler case is the temporally independent experiments, where $\mathbf{z_{t_{1}}}\perp\mathbf{z}_{t_{2}}$ for $t_{1},t_{2}\in[T]$, $t_{1}<t_{2}$. In this scenario, we have [^6] We clarify that the $T$ in the upper right corner of the vector and matrix always represents transposition, in this paper.
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ (33)
@@
$\operatorname{Var}\left((\mathbf{z}-w\mathbf{1})^{T}B\mathbf{z}\right)=O(\frac{1}{n})$ (34)
@@
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t}\right)$ (36)
$+2\sum_{t_{1}<t_{2}}\operatorname{Cov}$ $[(\mathbf{z}_{t_{1}}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t_{1}},(\mathbf{z}_{t_{2}}-\frac{\sum_{k=1}^{T}c_{k}}{T}\mathbf{1})^{T}B\mathbf{z}_{t_{2}}]$
@@ 4 Further Intuition: Increase the Variation of Treatment Exposures
In line with the principles of statistical learning [^7] As will be demonstrated in the simulation study, merging experimental data with different treatment proportions may not only fail to improve traditional estimators, such as the HT estimator and the difference-in-means estimator but can even degrade their performance., successful model training relies on sufficient variation in the feature space. We note that increasing the variation in covariates $X$ helps reduce the variance of the OLS estimator, $\sigma^{2}(X^{T}X)^{-1}$, which represents the classical case where unbiasedness is achieved. However, as shown in Section 3, under regular interference, the primary challenge lies in the bias term. Generally, greater variation in data can also reduce the bias of the machine learning model. For instance, in the Ridge regression, increasing data variation can not only reduce variance but also decreases bias. Similarly, in kernel regression with a Gaussian kernel, clustered data points result in a less informative kernel matrix, where most entries approach 1. This limits model flexibility and increases bias.
@@
$\operatorname{trace}\left(B^{T}B\operatorname{Cov}[\mathbf{z}]\right)$ (39)
@@
$\operatorname{trace}\left(B^{T}B\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T})(\mathbf{z}_{t}-\frac{\sum_{k=1}^{T}c_{k}}{T})^{T}\right)\right)$ (40)
for $T$-step experiment. Regretfully, these terms are concerned with the elements of second order term $B^{T}B$ for specific interference matrix $B$ and the network topology $A$ behind, which complicates deriving useful conclusions, such as the monotonicity with respect to the number of steps merged, $T$. This contrasts with the discussion of variance, where we only need to establish conclusions about the order. Therefore, we position this section as providing informative intuition.
@@ 5.1 Basic Setting
We begin by outlining the setup of our simulation study. Rather than using a simple random graph, we utilize a Facebook social network FB-Stanford3111111The network topology data can be found in https://networkrepository.com/socfb-Stanford3.php [58], which has also been employed in semi-synthetic experiments in [65, 18]. The network topology $\mathcal{G}$ consists of $|\mathcal{V}|=11586$ nodes and $|\mathcal{E}|=568309$ edges, where each node represents a person, and the edges indicate friendships or social relationships. Based on this real network, we generate the potential outcomes with specific models in the subsequent sections to validate the effectiveness of our methodology.
@@ 5.5 Multihop interference
$\text{GATE}=(\beta_{0}+\beta_{1}+\frac{\beta_{2}}{n}\mathbf{1}^{T}B\mathbf{1})-\beta_{0}\approx 2.489$ (47)
@@ A.1 Proof of Theorem 1
$\hat{\theta}=(X(\mathbf{z})^{T}X(\mathbf{z}))^{-1}X(z)^{T}Y$ (52)
@@
$X^{T}X=\begin{pmatrix}n&\mathbf{z}^{T}\mathbf{1}\\ \mathbf{z}^{T}\mathbf{1}&\|\mathbf{z}\|_{2}^{2}\end{pmatrix}=\begin{pmatrix}n&d\\ d&d\end{pmatrix}$ (53)
@@
$(X^{T}X)^{-1}X^{T}=\frac{1}{d(n-d)}\begin{pmatrix}d(\mathbf{1}-\mathbf{z})^{T}\\ (n\mathbf{z}-d\mathbf{1})^{T}\end{pmatrix}$ (54)
@@
$\hat{\theta}$ $=(X^{T}X)^{-1}X^{T}Y$ (56)
$=\begin{pmatrix}I_{2}&(X^{T}X)^{-1}X^{T}B\mathbf{z}\end{pmatrix}\beta+(X^{T}X)^{-1}X^{T}\mathbf{\epsilon}$
$=\begin{pmatrix}I_{2}&\frac{1}{d(n-d)}\begin{pmatrix}d(\mathbf{1}-\mathbf{z})^{T}B\mathbf{z}\\ (n\mathbf{z}-d\mathbf{1})^{T}B\mathbf{z}\end{pmatrix}\end{pmatrix}+(X^{T}X)^{-1}X^{T}\mathbf{\epsilon}$
@@
$\hat{\theta}_{1}=\beta_{1}+\beta_{2}\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{T}B\mathbf{z}+\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{T}\epsilon$ (57)
@@
$\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\beta_{2}\frac{1}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B))=\frac{\beta_{2}}{nc(1-c)}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}]-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)$ (58)
Next, we evaluate the term $\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}]$, which is given by:
$\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}]$ $=\mathbb{E}\left[\operatorname{trace}\left((\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}\right)\right]$ (59)
$=\mathbb{E}\left[\operatorname{trace}\left(\mathbf{z}(\mathbf{z}-c\mathbf{1})^{T}B\right)\right]$
$=\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-c\mathbf{1})^{T}\right]B\right)$
$=\operatorname{trace}\left((c\frac{d-1}{n-1}-c^{2})\mathbf{1}\mathbf{1}^{T}B\right)$
$=-\frac{c(1-c)}{n-1}\operatorname{trace}(\mathbf{1}\mathbf{1}^{T}B)$
@@
$\operatorname{Cov}[(n\mathbf{z}-d\mathbf{1})^{T}B\mathbf{z},(n\mathbf{z}-d\mathbf{1})^{T}\mathbf{\epsilon}]=0$ (61)
@@
$\operatorname{Var}(\hat{\tau})=(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}\right)+\frac{1}{\left(nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{T}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (62)
@@
$\operatorname{Var}_{e}$ $=\frac{1}{(d(n-d))^{2}}\mathbb{E}[(n\mathbf{z}-d\mathbf{1})^{T}(n\mathbf{z}-d\mathbf{1})]\sigma^{2}_{e}$ (63)
$=\left(\frac{n}{d(n-d)}\right)^{2}\mathbb{E}[(\mathbf{z}-\frac{d}{n}\mathbf{1})^{T}(\mathbf{z}-\frac{d}{n}\mathbf{1})]\sigma^{2}_{e}$
@@
$(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{T}B\mathbf{z}\right)$ (64)

### v2

@@ Abstract
A/B tests on modern platforms, including social networks and two-sided marketplaces, often encounter challenges from network interference. Extant research on causal inference under network interference based on single experiments often incurs substantial bias, particularly with complex interference. In this paper, we demonstrate the statistical benefit of merging the data of multiple experiments with different treatment proportions. Sequential experimentation with increasing traffic is often known as the ramp-up process in tech companies and is commonly applied due to operational reasons such as risk management and cost control. Beyond operational considerations, we find that regression-based estimators trained with such merged data achieve substantial bias reduction even with simple randomization schemes. Since the ramp-up process is standard practice, our methodology does not require additional experiment resources, providing a true plug-and-play solution that is easy to implement. We present a rigorous theoretical analysis of the bias and variance for the linear regression estimator of the global average treatment effect (GATE) under general linear network interference. We show that bias plays a dominant role in practical settings, and we highlight the bias reduction achieved by merging data from experiments with different treatment proportions. We provide an intuitive explanation for the bias reduction and, through the same lens, illustrate the synergistic effect of pooling experiments and cluster-level randomization. Furthermore, we consider a more advanced estimator based on graph neural networks (GNN). Through extensive simulation studies across novel and challenging scenarios, we show that the regression-based estimator benefits remarkably from our methodology, achieving outstanding statistical performance.
@@ 1 Introduction
A/B tests, also known as randomized controlled trials, have been widely adopted by modern online platforms as a paradigm of data-driven decision-making to support rapid product iteration [30, 42]. In most cases, the causal estimand is the GATE, since the experimenter is concerned with how much improvement on performance metrics, such as user engagement, would be achieved by introducing a new product feature to all users versus maintaining the status quo. In complex systems like social networks and marketplaces, the classic Stable Unit Treatment Value Assumption (SUTVA) in causal inference is often violated in cases of social interaction and competition, where the outcome of one user is influenced by the treatments of neighboring users. In such scenarios, the challenge of causal inference under interference naturally arises. This interference, mediated through a graph—such as a social network or market bipartite—is called network interference.
@@ 2.1 Preliminary
We adopt the Neyman-Rubin potential outcome framework [59] and consider a finite population of $n$ units. The treatment vector is represented as $\mathbf{z}=\left(z_{1},z_{2},\ldots,z_{n}\right)\in\{0,1\}^{n}$. Without assuming the Stable Unit Treatment Value Assumption (SUTVA), we define the potential outcome of unit $i$ as $Y_{i}=Y_{i}(\mathbf{z})$ to account for its possible dependency on the treatments of other units.
@@ 2.2 General Linear Interference Model
Where $\beta_{0},\beta_{1},\beta_{2}$ and $B$ are all unknown parameters, and we assume $\beta_{2}\neq 0$, indicating that there is indeed an interference effect. This model can be written more compactly as:
@@
If we take $B=D^{-1}A$, then our potential outcome model is reduced to the classic linear-in-means model [2], formatted as follows:
@@
We mention that removing the degree normalization, i.e. setting $B=A$ results in another common case where the exposure condition shifts from the proportion to the number of treated neighbors, thereby characterizing a different class of social interaction.
@@
$B=(\mathbf{1}\mathbf{1}^{\top}-I_{n})\odot\left(\sum_{m=1}^{M}r_{m}(D^{-1}A)^{m}\right)$ (6)
@@ 2.4 Regression-based Estimator
$\hat{\tau}(f)=\frac{1}{n}\mathbf{1}^{\top}\left(f(\mathbf{1},A)-f(\mathbf{0},A)\right)$ (8)
@@ 3 Analysis on Linear Regression Estimator
In this section, we perform a detailed analysis of the simple linear regression model [^1] We refer to it as ”simple” because it does not incorporate knowledge of graph topology; that is, $f(\mathbf{z},A)=f(\mathbf{z})$ for $f\in\mathcal{F}_{lr}$. As will be demonstrated in the subsequent subsections, this regression model supports precise analysis on bias and variance and provides a clear illustration of the role of merging experimental data for training. Many of the phenomena observed with this simple regression model can also be generalized to other regression models, such as those in $\mathcal{F}_{gnn}$, which will be discussed in the simulation study. On the other hand, we also note that precise statistical analysis would become infeasible if a component of network exposure is involved in the regression model. In summary, the linear regression model examined here is both straightforward and sufficiently powerful for illustration, allowing for the development of insightful conclusions. To be specific, the considered linear regression model is given by:
$Y=X(\mathbf{z})\theta+\mathbf{\epsilon_{\theta}}+\epsilon^{\prime}$ (9)
here $X(\mathbf{z})=(\mathbf{1},\mathbf{z})$, $\theta=(\theta_{0},\theta_{1})^{\top}$, and $\epsilon^{\prime}$ is the error term.
@@
$\hat{\theta}=(X(\mathbf{z})^{\top}X(\mathbf{z}))^{-1}X(\mathbf{z})^{\top}Y$ (11)
@@ 3.1 One-step Experiment
$\operatorname{Bias}(\hat{\tau})$ $=\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B))$ (12)
$=\frac{\beta_{2}}{nc(1-c)}\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-c\mathbf{1})^{\top}\right]B\right)-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$
@@
$\operatorname{Var}(\hat{\tau})/\beta_{2}^{2}$ $=(\frac{1}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)+\frac{1}{\left(\beta_{2}nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (13)
@@
$\mathbf{1}^{\top}D^{-1}A^{2}D^{-1}\mathbf{1}=O(n)$ (21)
For instance, in a ring graph, we have $\sum_{j}(\sum_{i}B_{ij})=\Theta(n)$, while in a star graph—where one opinion leader connects to all other units—we have $\sum_{j}(\sum_{i}B_{ij})=\Theta(n^{2})$. In the context of Example 1, a sufficient condition for satisfying this assumption is restricting the growth rate to a constant $\kappa$, a condition commonly employed in related works [64]. However, a restricted growth rate is not a necessary condition. As a counterexample, consider the complete graph, where $\mathbf{1}^{\top}D^{-1}A^{2}D^{-1}\mathbf{1}=n$; despite being the densest graph, it lacks any prominent opinion leaders. We further assert that this assumption is reasonable for typical social networks like WeChat, where, even with over a billion nodes, it is unlikely for any single unit to have more than 100,000 friends while maintaining significant influence over each one of them.
@@ 3.3 $T$-step Experiment
By Cauchy-Schwarz inequality and basic algebra, we can verify that
@@
In addition, we can apply the Corollary 2 in the reverse direction: given the proportions, $(c_{t})_{t=1}^{T}$, removing the data with treatment proportion $c_{t}$ from the merged data set could potentially reduce bias further. Repeating this procedure suggests the final solution would involve merging only the experimental datasets with proportions $c_{1}$ and $c_{T}$. This outcome underscores the particular significance of the $T=2$ case, and we will further verify it in the simulation study. Nevertheless, merging multistep experimental data can be beneficial under nonlinear interference, as discussed in Section 5. There, we will also outline the recommended practice for applying our methodology, in conjunction with the results of the simulation study.
@@
The simpler case is the temporally independent experiments, where $\mathbf{z_{t_{1}}}\Perp\mathbf{z}_{t_{2}}$ for $t_{1},t_{2}\in[T]$, $t_{1}<t_{2}$. In this scenario, we have:
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ (33)
@@
$\operatorname{Var}\left((\mathbf{z}-w\mathbf{1})^{\top}B\mathbf{z}\right)=O(\frac{1}{n})$ (34)
@@
$\operatorname{Var}\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ $=\sum_{t=1}^{T}\operatorname{Var}\left((\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t}\right)$ (36)
$+2\sum_{t_{1}<t_{2}}\operatorname{Cov}$ $[(\mathbf{z}_{t_{1}}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t_{1}},(\mathbf{z}_{t_{2}}-\frac{\sum_{k=1}^{\top}c_{k}}{T}\mathbf{1})^{\top}B\mathbf{z}_{t_{2}}]$
@@ 4 Further Intuition: Increase the Variation of Treatment Exposures
In line with the principles of statistical learning [^5] As will be demonstrated in the simulation study, merging experimental data with different treatment proportions may not only fail to improve traditional estimators, such as the HT estimator and the difference-in-means estimator but can even degrade their performance., successful model training relies on sufficient variation in the feature space. We note that increasing the variation in covariates $X$ helps reduce the variance of the OLS estimator, $\sigma^{2}(X^{\top}X)^{-1}$, which represents the classical case where unbiasedness is achieved. However, as shown in Section 3, under regular interference, the primary challenge lies in the bias term. Generally, greater variation in data can also reduce the bias of the machine learning model. For instance, in the Ridge regression, increasing data variation can not only reduce variance but also decreases bias. Similarly, in kernel regression with a Gaussian kernel, clustered data points result in a less informative kernel matrix, where most entries approach 1. This limits model flexibility and increases bias.
@@
$\operatorname{trace}\left(B^{\top}B\operatorname{Cov}[\mathbf{z}]\right)$ (39)
@@
$\operatorname{trace}\left(B^{\top}B\left(\sum_{t=1}^{T}(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T})(\mathbf{z}_{t}-\frac{\sum_{k=1}^{\top}c_{k}}{T})^{\top}\right)\right)$ (40)
for $T$-step experiment. Regretfully, these terms are concerned with the elements of second order term $B^{\top}B$ for specific interference matrix $B$ and the network topology $A$ behind, which complicates deriving useful conclusions, such as the monotonicity with respect to the number of steps merged, $T$. This contrasts with the discussion of variance, where we only need to establish conclusions about the order. Therefore, we position this section as providing informative intuition.
@@ 5.1 Basic Setting
We begin by outlining the setup of our simulation study. Rather than using a simple random graph, we utilize a Facebook social network FB-Stanford3999The network topology data can be found in https://networkrepository.com/socfb-Stanford3.php [58], which has also been employed in semi-synthetic experiments in [65, 18]. The network topology $\mathcal{G}$ consists of $|\mathcal{V}|=11586$ nodes and $|\mathcal{E}|=568309$ edges, where each node represents a person, and the edges indicate friendships or social relationships. Based on this real network, we generate the potential outcomes with specific models in the subsequent sections to validate the effectiveness of our methodology.
@@ 5.5 Multihop interference
$\text{GATE}=(\beta_{0}+\beta_{1}+\frac{\beta_{2}}{n}\mathbf{1}^{\top}B\mathbf{1})-\beta_{0}\approx 2.489$ (47)
@@ A.1 Proof of Theorem 1
$\hat{\theta}=(X(\mathbf{z})^{\top}X(\mathbf{z}))^{-1}X(z)^{\top}Y$ (52)
@@
$X^{\top}X=\begin{pmatrix}n&\mathbf{z}^{\top}\mathbf{1}\\ \mathbf{z}^{\top}\mathbf{1}&\|\mathbf{z}\|_{2}^{2}\end{pmatrix}=\begin{pmatrix}n&d\\ d&d\end{pmatrix}$ (53)
@@
$(X^{\top}X)^{-1}X^{\top}=\frac{1}{d(n-d)}\begin{pmatrix}d(\mathbf{1}-\mathbf{z})^{\top}\\ (n\mathbf{z}-d\mathbf{1})^{\top}\end{pmatrix}$ (54)
@@
$\hat{\theta}$ $=(X^{\top}X)^{-1}X^{\top}Y$ (56)
$=\begin{pmatrix}I_{2}&(X^{\top}X)^{-1}X^{\top}B\mathbf{z}\end{pmatrix}\beta+(X^{\top}X)^{-1}X^{\top}\mathbf{\epsilon}$
$=\begin{pmatrix}I_{2}&\frac{1}{d(n-d)}\begin{pmatrix}d(\mathbf{1}-\mathbf{z})^{\top}B\mathbf{z}\\ (n\mathbf{z}-d\mathbf{1})^{\top}B\mathbf{z}\end{pmatrix}\end{pmatrix}+(X^{\top}X)^{-1}X^{\top}\mathbf{\epsilon}$
@@
$\hat{\theta}_{1}=\beta_{1}+\beta_{2}\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}B\mathbf{z}+\frac{1}{d(n-d)}(n\mathbf{z}-d\mathbf{1})^{\top}\epsilon$ (57)
@@
$\mathbb{E}[\hat{\theta}_{1}]-(\beta_{1}+\beta_{2}\frac{1}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B))=\frac{\beta_{2}}{nc(1-c)}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}]-\frac{\beta_{2}}{n}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$ (58)
Next, we evaluate the term $\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}]$, which is given by:
$\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}]$ $=\mathbb{E}\left[\operatorname{trace}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)\right]$ (59)
$=\mathbb{E}\left[\operatorname{trace}\left(\mathbf{z}(\mathbf{z}-c\mathbf{1})^{\top}B\right)\right]$
$=\operatorname{trace}\left(\mathbb{E}\left[\mathbf{z}(\mathbf{z}-c\mathbf{1})^{\top}\right]B\right)$
$=\operatorname{trace}\left((c\frac{d-1}{n-1}-c^{2})\mathbf{1}\mathbf{1}^{\top}B\right)$
$=-\frac{c(1-c)}{n-1}\operatorname{trace}(\mathbf{1}\mathbf{1}^{\top}B)$
@@
$\operatorname{Cov}[(n\mathbf{z}-d\mathbf{1})^{\top}B\mathbf{z},(n\mathbf{z}-d\mathbf{1})^{\top}\mathbf{\epsilon}]=0$ (61)
@@
$\operatorname{Var}(\hat{\tau})=(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)+\frac{1}{\left(nc(1-c)\right)^{2}}\mathbb{E}[(\mathbf{z}-c\mathbf{1})^{\top}(\mathbf{z}-c\mathbf{1})]\sigma^{2}_{e}$ (62)
@@
$\operatorname{Var}_{e}$ $=\frac{1}{(d(n-d))^{2}}\mathbb{E}[(n\mathbf{z}-d\mathbf{1})^{\top}(n\mathbf{z}-d\mathbf{1})]\sigma^{2}_{e}$ (63)
$=\left(\frac{n}{d(n-d)}\right)^{2}\mathbb{E}[(\mathbf{z}-\frac{d}{n}\mathbf{1})^{\top}(\mathbf{z}-\frac{d}{n}\mathbf{1})]\sigma^{2}_{e}$
@@
$(\frac{\beta_{2}}{nc(1-c)})^{2}\operatorname{Var}\left((\mathbf{z}-c\mathbf{1})^{\top}B\mathbf{z}\right)$ (64)